OpenStudy (anonymous):

simple exponential equation help! solve (without using logs) 3(^2x+1)=70

OpenStudy (dumbcow):

i get x = 1.43 but i used logs so it doesnt count

OpenStudy (dumbcow):

why cant you use logs

OpenStudy (anonymous):

how do you do it with logs then?

OpenStudy (anonymous):

haha i was already excited what the solution is xD

OpenStudy (anonymous):

well it's grade 11 functions and apps, and logs is apparently advanced functions so she doesnt want us using them. but whatever, i'll use logs.

OpenStudy (dumbcow):

take log base 3 of both sides ->2x+1 = log3(70) get x by itself _>2x = log3(70) -1 ->x = (1/2)*(log3(70)-1) use change of base formula to find log3(70)

OpenStudy (dumbcow):

log3(70) = log(70)/log(3) sorry you prob didnt know what i meant

OpenStudy (dumbcow):

wait sorry i think i know what your teacher wanted 3^2x+1 = 3^2x*3^1 = 70 divide 3 on both sides 3^2x = 70/3 square root both sides 3^x = sqrt(70/3) = 4.83 then estimate your best guess for x